CURRICULUM VITAE

GAVRIL FARKAS

Professor of Mathematics Humboldt Universität zu Berlin Institut für Mathematik, Unter den Linden 6 10099 Berlin, Germany Phone: 49 (30) 2093-5412 e-mail: farkas@math.hu-berlin.de web: http://www-irm.mathematik.hu-berlin.de/˜ farkas/

Personal Information

Date of birth: March 19, 1973 Place of birth: Nagyvárad (Oradea), Romania Citizenship: Hungary, Romania, Germany

Area of specialization

Algebraic geometry with connection to other fields especially topology, combinatorics and group theory

Positions

2007–: Humboldt Universität zu Berlin, Full Professor (W3) 2006–2007: University of Texas at Austin, Associate Professor 2004-2006: University of Texas at Austin, Assistant Professor (tenured 2006) 2003–2004: Princeton University, Assistant Professor 2000–2003: University of Michigan, Ann Arbor, Postdoc

Education

1996–2000: Universiteit van Amsterdam, PhD June 2000

Thesis: The birational geometry of the moduli space of curves. Adviser: G. van der Geer 1995–1996: Mathematical Research Institute in the Netherlands, MSc in Mathematics 1991–1995: “Babes-Bolyai” University, Kolozsvár (Cluj), Romania, Bachelor of Science in Mathematics, specialization: geometry and topology

Other professional experience

• Harvard University: invited visitor, Sep-Dec 1998, Oct-Dec 1999

• Newton Institute, Cambridge UK: invited visitor, June-July 2002, Feb-April 2011 • Korean Institute for Advanced Studies: invited visitor, March-Apr 2005, Sep 2009 • Mittag-Leffler Institute Stockholm: Visiting Professor, April-June 2007, October-

November 2021

• IMPA, Rio de Janeiro: Visiting Professor, January-February 2007

• MSRI Berkeley: Research Member, February-April 2009

• University of Georgia, Athens: Visiting Professor, March-April 2010 • Institut Henri Poincare, Paris: Visiting Professor, April-May 2010

• Instituto Nazionale di Alta Matematica, Rome: Visiting Professor, Oct-Dec 2011

• Vietnamese Institute of Advanced Studies Hanoi: Visiting Professor, Feb 2014

• Warwick University: Visiting Professor, Feb-March 2015

• Università di Roma Tre: Visiting Professor, October-December 2015

• MSRI Berkeley: Research Member, February-April 2019

• Rényi Institute of Mathematics Budapest: Visiting Professor, March-August 2022

Awards, prizes

• ERC Advanced Grant SYZYGY (2019-2025)

• Full member of the Berlin-Brandenburg Akademie der Wissenschaften 2022

• First Math+ Distinguished Fellow 2022

• Ad Astra Prize for Excellence in Scientific Research 2014

• Alfred P. Sloan Research Fellowship (2005-2008)

Grants and fellowship

• DFG Project Syzygien und Moduli (2019-2022)

• Principal Investigator in the Math+ Excellence Center in Mathematics (2019-2026)

• Project leader in the DFG Sonderforschungsbereich 647 Raum-Zeit-Materie (2009-2012 and 2013-2016)

• Principal Investigator in the DFG International Research Training Group 1800 Moduli spaces and automorphic forms Berlin-Amsterdam-Leiden (2012-2016)

• Principal Investigator in the DFG Schwerpunkt 1489 Algorithmische und Experimentelle Methoden in Algebra, Geometrie und Zahlentheorie (2010-2013 and 2013-2016)

• National Science Foundation, research grant DMS-0500747 (2005-2008)

• National Science Foundation, research grant DMS-0140520 (2002-2005)

• Principal Investigator on the NSF Research Training Grant at the University of Texas at Austin (with D. Ben-Zvi, D. Freed, S. Keel and K. Uhlenbeck)

• Texas Summer Research Assignment 2006, Project: Geometry of moduli spaces

• Rackham Faculty Fellowship, University of Michigan, June-August 2001

• Special prize of the Romanian Mathematical Society at the Romanian Mathematics Olympiad, 1991

• Winner of the “Traian Lalescu” Mathematics Contest for Romanian undergraduate students in 1992 and 1993

Editorial work

• 2013–: Managing Editor (with Y. Tschinkel) of the journal Algebraic Geometry

• 2016–: Editor of Algebra and Number Theory

• 2008–: Editor of Documenta Mathematica

• 2016–: Editor of the De Gruyter Series Studies in Mathematics

• 2022–: Editor of EMS Monographs in Mathematics

• 2019–: Editor of EMS Surveys in Mathematical Sciences

• 2008–: Editor of Mathematica (Cluj), the Journal of the Romanian Academy

Volumes edited

• Editor (with I. Morrison) of the Handbook of Moduli, 3 volumes, Intl. Press 2013

• Editor (with C. Faber and G. van der Geer) of the volume K3 surfaces and their moduli, Progress in Mathematics vol. 315, Birkhäuser 2016

• Editor (with C. Faber and R. de Jong) of the volume Geometry and Arithmetic, Series of Congress Reports 2012, European Mathematical Society

• Editor (with. G. van der Geer, M. Shen and L. Taelman) of the volume Rationality of varieties, Progress in Mathematics vol. 342, Birkhäuser 2021

PhD Students

(1) P. Larsen, PhD 2010-Applied Mori theory of the moduli space of rational pointed curves (2) N. Tarasca, PhD 2012-Geometric cycles in the moduli space of curves

(3) M. Lelli-Chiesa, PhD 2012-Gieseker-Petri divisors and Brill-Noether theory of K3 sections-Mario Baldassari prize of the UMI

(4) F. Müller, PhD 2013-Models of moduli spaces of curves of small genus

(5) G. Bruns, PhD 2016-Divisors on moduli spaces of level curves

(6) A. Peterson, PhD 2015 at Universiteit van Amsterdam advised jointly with G. van der Geer-Divisors on moduli spaces of polarized K3 surfaces

(7) E. Sertöz, PhD 2017-Enumerative geometry of double spin curves

(8) I. Kadiköylu, PhD 2018, Rank stratifications of spaces of quadrics and moduli of curves (9) D. Agostini, PhD 2018-Syzygies of algebraic varieties and moduli

(10) I. Barros Reyes, PhD 2018-K3 surfaces and strata of holomorphic differentials

(11) M. Ungureanu, PhD 2018-Enumerative formulas of De Jonquieres type on curves

(12) I. Schwarz, PhD 2020-Birational geometry of moduli spaces of pointed curves

(13) C. Maestro Perez, PhD 2021-Aspects of the geometry of Prym varieties and their moduli (14) A. Bud, PhD 2022-Moduli of Prym curves and strata of differentials

Postdoctoral fellows

(1) 2005-2007: Deepak Khosla (PhD 2005, Harvard University with J. Harris) Research mentoring at the University of Texas in Austin

(2) 2008-2012: Cristina Manolache (PhD 2008, SISSA Trieste with B. Fantechi) SFB 647 at the HU Berlin; currently Research Fellow at Imperial College

(3) 2009-2010: Michele Bolognesi (PhD 2006, Université de Nice with C. Pauly) Assistant position at HU Berlin; currently Prof. at U. Montpellier

(4) 2010-2013: Paul Larsen (PhD 2010, Humboldt Universität zu Berlin)

Research mentoring, assistant position at HU Berlin

(5) 2012-2013, 2014-2018: Frank Gounelas (PhD 2012, Oxford with V. Flynn)

Financed by the SFB 647 and HU Berlin; currently Substitute Prof. at Uni. Göttingen

(6) 2013-2016: Fabio Tonini (PhD 2013, Scuola Normale Sup. di Pisa with A. Vistoli)

(7) 2013-2016: Michael Kemeny (PhD 2015, Universität Bonn with D. Huybrechts) Research mentoring (SPP 1489), currently Assoc. Prof. at U. Wisconsin

(8) 2013-2015: Jarod Alper (PhD 2008, Stanford Univ. with R. Vakil)

Humboldt Forschungsstipendium; currently Assoc. Prof. at U. Washington (9) 2013-2014: Alwaleed Kamel (PhD 2010, Saitama University with F. Sakai)

Georg Foster Fellowship; currently Prof. at Sohang U.-Egypt.

(10) 2014-2016: Benjamin Bakker (PhD 2011, Princeton Univ. with R. Pandharipande) Financed by the IRTG 1800; currently Assoc. Prof. at U. Illinois Chicago

(11) 2014-2016: Jérémy Guéré (PhD 20105, Université Paris 6 with A. Chiodo) Einstein Foundation Berlin; currently Maitre de Conf. at U. Grenoble

(12) 2015-2016: Valentin Tonita (PhD 2011, UC Berkeley with A. Givental) Financed by the SFB 647

(13) 2016-2017: Olaf Bergvall (PhD 2016, KTH Stockholm with C. Faber)

Einstein Foundation Berlin, currently Assoc. Prof. at Mälardalen U.

(14) 2017-2020: Jason van Zelm (PhD 2018, University of Liverpool with N. Pagani) Position financed by the Einstein Foundation Berlin

(15) 2019-2022: Andrey Soldatenkov (PhD 2014, U. Moscow with M. Verbitsky) Position financed by the DFG, currently permanent researcher at Steklow Inst.

(16) 2018-2022: Daniele Agostini (PhD 2018, HU Berlin)

Position financed by HU Berlin, currently tenure-track W3 Professor at U. Tübingen (17) 2020-2022: András Lőrincz (PhD 2016, U. Connecticut with J. Weyman)

Dirichlet Postdoc of the Berlin Mathematical School, currently tenure-track Ass. Professor at U. Oklahoma

(18) 2020-2023: Zhuang He (PhD 2020, Northeastern U. with A.-M. Castravet) Position financed by ERC

(19) 2020-2023: Carl Lian (PhD 2020, Columbia U. with J. de Jong)

NSF Postdoc

(20) 2021-2024: Ruijie Yang (PhD 2021, Stony Brook U. with R. Lazarsfeld) Position financed by ERC

(21) 2022-2025: Andres Rojas (PhD 2021, U. Barcelona with M. Lahoz and J.-C. Naranjo) Position financed by ERC

(22) 2022-2025: Jieao Song (PhD 2022, U. Sorbonne with O. Debarre)

Dirichlet Postdoc of the Berlin Math. School

(23) 2021-2027: Andrea di Lorenzo (PhD 2020, Scuola Norm. Sup. Pisa with A. Vistoli) Position financed by HU Berlin

Professional activities

Organization of numerous conferences and workshops, among them:

• Resonance, topological invariants of groups, moduli, November 2022, HU Berlin

• Varieties, polyhedra, algorithms, Thematic Einstein Semester in Algebraic Geometry organised by the Math+ Excellence Center, September 2019-February 2020, with P. Bürgisser and C. Haase

• Brill-Noether theory, November 2019, HU Berlin

• Moduli in Budapest, April 2019, Rényi Institute of Mathematics Budapest, with A. Némethi, R. Pandharipande and A. Stipsicz

• Rationality of algebraic varieties, April 2019, Schiermonnikoog, with G. van der Geer, M. Shen and L. Taelman

• Higgs bundles and K3 surfaces, July 2017, HU Berlin, with H. Esnault

• K3 surfaces, sheaves and moduli, February 2017, HU Berlin with R. Pandharipande • Satellite Conference of the European Congress of Mathematics, July 2016, with K.

Hulek, S. Müller-Stach and M. Schütt

• Oberwolfach Workshop Recent trends in algebraic geometry, June 2023 with O. Debarre, Y. Tschinkel and R. Vakil

• Oberwolfach Workshop Classical Algebraic Geometry, June 2014, June 2016, July 2018 and June 2021 with O. Debarre, D. Eisenbud and R. Vakil

• Geometry of algebraic varieties, October 2015, with A. Ortega, conference on the occasion of A. Verra’s 65th birthday

• Recent developments in algebraic and arithmetic geometry, August 2015, Siena, with L. Chiantini, G. van der Geer and J. Kramer

• Motivic invariants associated to K3 and abelian geometries, February 2015, HU Berlin, with R. Pandharipande

• Recent developments in algebraic and arithmetic geometry, August 2014, Budapest, Rényi Institute of Mathematics

• K3 surfaces and their moduli, May 2014, Schiermonnikoog, with C. Faber and G. van der Geer

• Oberwolfach Workshop Moduli spaces in algebraic geometry, January 2010 and February 2013, with D. Abramovich, L. Caporaso and S. Kebekus

• Birational geometry and stability of moduli stacks, January-February 2014, with I. Morrison, workshop at the Vietnamese Institute for Advanced Studies, Hanoi

• Geometry and topology of moduli, Berlin, Humboldt Universität, October 2012

• Banff BIRS Workshop Syzygies in algebraic geometry with the exploration of a connection to string theory August 2012, with L. Ein, D. Eisenbud and I. Peeva

• Geometry and Arithmetic (Schiermonnikoog 2010), on the occasion of Gerard van der Geer’s 60th birthday, with C. Faber and R. de Jong

• Complex Algebraic Geometry, Berlin, Humboldt Universität, October 2009, on the occasion of H. Kurke’s 70th birthday, with K. Altmann

• Moduli, Berlin, Humboldt Universität, August 2009, with J. Kramer

• Romanian-German Conference in Algebraic Geometry, Sibiu/Hermannstadt, May 2009, with M. Aprodu

• Warwick Algebraic Geometry Symposium (2007-2008), with A. Corti, S. Mukai and M. Reid

• Co-organizer of the Northern German Algebraic Geometry Seminar, as of 2007

• Organizer of the Princeton Algebraic Geometry Seminar (2003-2004)

• Co-organizer of the Texas Algebraic Geometry Seminar (2005-2007)

• Curves and their moduli, University of Michigan, Ann Arbor, March 2002, with A. Gibney and T. Nevins

Service

• Director on the Institute of Mathematics of the Humboldt Univ. Berlin (2021-)

• Cochair of the Berlin Matematical School (2023-)

• Principal Investigator and member of the Council of the Excellence Center in Mathematics Math+.

• Member of the Executive Board of the Berlin Mathematical School

• Member of the Jury for the annual Euler Vorlesung at the Neues Palais in Potsdam (as of 2015)

• I have reviewed numerous times proposals submitted to the national funding agencies in USA (NSF), ERC, Germany (DFG), Chile, Italy, Norway, Poland, Portugal, Romania, Russia, Sweden, Switzerland and The Netherlands

• Senior speaker at GAEL 2009 (Géométrie Algébrique en Liberté), Leiden, June 2009 • Invited series of lectures at numerous schools including Cologne (August 2006),

Lisbon (September 2011), Daejeon (January 2007 and March 2013), Rio de Janeiro (February 2007), Hamburg (August 2013), Zacatecas (June 2006), Nordfjordeid (June 2012), Bedlewo (July 2010), Grenoble (June 2011), Stony Brook (July 2014), Guanajuato (March 2016), Helvetic Algebraic Geometry Seminar (Konolfingen 2018)

• Mentor “Warm-Up Workshop for AMS Summer Institute in Algebraic Geometry”, Seattle, July 2005 and “Algebraic Geometry: Presentations by Young Researchers” Snowbird, Utah, July 2004

Selected talks

• “Complex geometry and geometric group theory”, Karlsruhe Institute of Technology, September 2022

• “Algebraic Geometry in Hannover”, on the occasion of K. Hulek’s 70th birthday, Hannover, September 2022

• Plenary Speaker at the Annual Meeting of the Portuguese Mathematical Society, Tomar, July 2022

• “Vector Bundles on Algebraic Curves 2022-New Trends”, on the occasion of P. Newstead’s 80th birthday, Warwick, July 2022

• “Algebraic Geometry in Roma Tre”, on the occasion of A. Verra’s 70th birthday, Rome, June 2022

• “Moduli of curves in Stockholm”, on the occasion of C. Faber’s 60th birthday, Mittag-Leffler Institute, June 2022

• “Géométrie Algebrique en l’honneur de Claire Voisin”, a conference to celebrate C. Voisin’s 60th birthday, Paris, May 2022

• Fourth annual conference of the SFB 195, Kaiserslautern, September 2021

• “Moduli spaces and modular forms”, Oberwolfach, January 2021

• “Hyperplane arrangements and the Japanese-Australian workshop on singularities”, Tokyo, December 2019

• “The geometry of algebraic varieties” a conference to celebrate O. Debarre’s 60th birthday, Luminy, September 2019

• “Facets of algebraic geometry”, a conference to celebrate W. Fulton’s 80th birthday, Ann Arbor, October 2019

• “Birational geometry and moduli”, MSRI Berkeley, May 2019

• “Discrete groups and moduli”, Nagoya, June 2019

• “Flat surfaces and algebraic curves”, Oberwolfach, September 2018

• “Moduli spaces”, ICM Satellite Conference, Campinas, July 2018

• “Topology and geometry”, in the memory of S. Papadima, Bucharest, May 2018

• “Automorphic forms and algebraic geometry”, Steklov Institute, St. Petersburg, May 2018

• “French-German meeting on complex geometry”, CIRM Luminy, April 2018 • “Complex algebraic geometry”, Barcelona, a conference to celebrate P. Pirola’s

60th birthday, February 2018

• “Algebraic geometry”, to celebrate A. Beauville’s 70th birthday, Beijing, Chinese Academy, October 2017

• “Instruments of algebraic geometry”, Institute of Mathematics of the Romanian Academy, September 2017

• “A view towards algebraic geometry”, Martha’s Vineyard, a conference to celebrate D. Eisenbud’s 70th birthday, May 2017

• “Cycles and moduli“ ETH Zürich, March 2017

• “Stability and moduli spaces”, American Institute of Mathematics, Jan. 2017

• “Algebraic geometry”, College de France, October 2016

• “Hyperkähler manifolds“, Milano, October 2016

• “British Algebraic Geometry Seminar“, Edinburgh, April 2016

• “Moduli and modular forms”, Oberwolfach Workshop, April 2016

• “Emmy Noether Lectures” (2 lectures), Bar Ilan University, January 2016

• “Moduli spaces in geometry” CIRM Luminy, October 2015

• “AMS Summer Institute in Algebraic Geometry”, Salt Lake City, July 2015

• “Mathematische Arbeitstagung”, Bonn, June 2015

• “Applications of syzygies”, Saarbrücken, July 2015, a conference to celebrate F.-O. Schreyer’s 60th birthday

• “Minimal resolutions, Betti numbers and combinatorics”, Edinburgh, May 2015 • “In onore di Guido Castelnuovo”, Venice, Instituto Veneto di Arte, Letteri e Scienzi,

May 2015

• “Geometry over non-closed fields”, Simons Foundation, Puerto Rico, March 2015 • “Tropical geometry in Europe”, TU Berlin, March 2015

• “Algebraic geometry”, Chinese Academy, Beijing, October 2014

• Joint meeting of the German and Polish Mathematical Societies, Poznan, September 2014

• “Algebraic geometry”, ICM satellite conference, Daejon, August 2014

• “Explicit moduli spaces and cryptography”, Rennes, June 2014

• “Irreducible symplectic varieties and moduli”, Lille, March 2014

• “Brill-Noether methods”, Bonn, Hausdorff Center, March 2014

• “Weihnachtskonferenz Geometrie und Zahlentheorie”, Karlsruhe, December 2013 • “AGNES, Boston, October 2013

• “Integrable systems and moduli spaces”, Banff, August 2013

• “Moduli spaces in algebraic geometry and physics” (3 lectures), Hamburg, August 2013

• “Algebraic Geometry”, Amsterdam, July 2013

• “Developments in moduli theory”, Kyoto, June 2013, a conference to celebrate S. Mukai’s 60th birthday

• “Tropical geometry in Europe”, Geneve, May 2013

• “Syzygies and moduli” (3 lectures), KAIST Daejon, Korea, March 2013

• “Representation theory, homological algebra and free resolutions”, MSRI Berkeley, February 2013

• “Log minimal model program for moduli spaces”, American Institute for Mathematics, Palo Alto, December 2012 • “Texas Geometry and Topology”, Houston, Texas, November 2012

• “Algebraic and complex geometry”, Hannover, September 2012, a conference on the occasion of Klaus Hulek’s 60th birthday

• “Komplexe Analysis”, Oberwolfach, August 2012

• “Classical Algebraic Geometry”, Mathematisches Forschungsinstitut Oberwolfach, June 2012

• “Georgia Algebraic Geometry Symposium”, Athens, Georgia, May 2012

• GAEL XX+ Grenoble, June 2012, to celebrate GAEL’s 20th anniversary

• “Geometry of moduli spaces”, Sophus Lie Center, Nordfjordeid, Norway, June 2012, a series of lectures together with M. Reinecke and C. Voisin

• “Geometria delle Varieta Algebriche Turin, March 2012, a conference to celebrate A. Conte’s 70th birthday

• “Modular forms and moduli spaces”, CIRM Luminy, October 2011

• “Periods and moduli spaces”, Korean Institute of Advanced Studies, Seoul, September 2011

• “Topics in moduli theory” (4 lectures), Instituto Superior Tecnico, Lisbon, September 2011

• “A celebration of algebraic geometry”, Cambridge Massachusetts August 2011, a conference in honor of J. Harris’ 60th birthday

• “Moduli of Riemann surfaces”, Park City, Utah, July 2011

• The Seventh Congress of Romanian Mathematicians, Brasov, July 2011

• “Vector bundles on algebraic curves”, Isaac Newton Institute Cambridge, June 2011

• “Moduli of curves and Gromov-Witten theory”, Grenoble June-July 2011

• “Moduli spaces and moduli stacks”, New York, Columbia University, May 2011 • “Groups and Algebraic Geometry”, Erlangen, May 2011-a conference on the ocas-

sion of the retirement of W. Barth

• Cambridge-Oxford-Warwick Seminar, Cambridge March 2011

• “Géometrie algèbrique complexe”, Luminy, January 2011

• “Classical Algebraic geometry”, Instituto Nazionale di Alta Matematica, Rome, February 2011

• Kinosaki Algebraic Geometry Symposium, Kinosaki (Japan), October 2010

• V Iberoamerican Congress on Geometry, Pucon (Chile), December 2010

• “Perspectives on algebraic varieties”, Levico Terme, a conference on the ocassion of the 60th birthday of F. Catanese, C. Ciliberto and A. Verra, September 2010

• “Moduli and birational geometry”, Pohang (Korea), August 2010

• IMPANGA Summer School on Algebraic Geometry, Bedlewo (Poland), July 2010 • “Classical Algebraic Geometry”, Mathematisches Forschungsinstitut Oberwol-

fach, June 2010

• Harvard-MIT Algebraic Geometry Seminar, March 2010.

• “After Carnival: An Algebraic Geometry Party”, Turin, February 2010.

• “Complex Algebraic geometry”, Seoul, September 2009.

• “Complex Algebraic geometry”, Mathematisches Forschungsinstitut Oberwolfach, August 2009

• “Classification of algebraic varieties” Schiermonnikoog, The Netherlands, May • MSRI Main Seminar, Berkeley March 2009

• Invited plenary speaker at the annual meeting of the German Mathematical Society, Erlangen, September 2008

• “Algebraic stacks” Trento, Italy, September 2008

• “Komplexe Analysis”, Mathematisches Forschungsinstitut Oberwolfach, August 2008

• “Algebraic Geometry”, Institutul de Matematica al Academiei Romane, Bucharest, July 2008

• “Algebraic Geometry”, a satellite conference of the European Congress of Mathematicians, Leiden, July 2008

• “Aspects of moduli”, Pisa, Italy, June 2008

• “Classical Algebraic Geometry”, Oberwolfach, June 2008

• Princeton Algebraic Geometry Seminar, March 2008

• “Recent progress on moduli of curves”, Banff, Canada, March 2008

• “Moduli spaces of geometric and algebraic structures”, Bonn, January 2008

• “Algebraic geometry and commutative algebra”, Tokyo, December 2007

• “Complex Algebraic Geometry”, Oberwolfach, October 2007

• “Modular forms and moduli spaces”, St. Petersburg, Russia, July 2007

• “Linear series and subschemes”, Ghent, Belgium, April 2007

• “Moduli spaces” Mittag-Leffler Institute Djursholm, May 2007

• “Moduli of curves” (7 lectures), IMPA Rio de Janeiro, February 2007

• Winter School on Algebraic Curves, National Institute for Mathematical Sciences, Daejeon (Korea)-series of lectures, January 2007

• “Algebraic Geometry and its Applications, Sogang Univ. Seoul, January 2007

• Workshop on Vector Bundles, Clay Math. Institute, Cambridge, October 2006 • “Meetkunde in de Herfsttij”, Leiden, The Netherlands, September 2006

• “Moduli spaces and combinatorics”, Banff, July 2006

• Cambridge-Oxford-Warwick Seminar, Oxford, June 2006.

• Invited Speaker, “Tópicos de geometria algebraica y temas afines” (4 lectures) Zacatecas, Mexico, June 2006

• “Classical Algebraic Geometry”, Oberwolfach, June 2006

• Algebraic Geometry Seminar, University of Chicago, January 2006

• Michigan-Ohio State Algebraic Geometry Meeting, Columbus, April 2006

• “Recent Trends in Higher Dimensional Algebraic Geometry”, Banff, April 2006 • AMS Summer Institute in Algebraic Geometry, Seattle, July 2005

• AMS/Taiwan Joint International Meeting, Taichung, Taiwan, December 2005 • Berkeley Algebraic Geometry Seminar, December 2005

• “Higher Dimensional Algebraic Varieties”, KIAS Workshop, Seoul, March 2005 • “Western Algebraic Geometry Seminar”, Seattle, March 2005

• “Texas Algebraic Geometry” Conference, Houston, May 2005

• “Cohomology of the Moduli Space of Curves” Conference, American Institute of Mathematics, March 2005

• “Compact Moduli Spaces”, American Institute of Mathematics, December 2004 • “Classical Algebraic Geometry”, Oberwolfach, July 2004

• European Conference on Gromow-Witten invariants, Trieste, June 2004

• Third Ibero-American Congress of Geometry, Salamanca, June 2004 • Harvard-MIT Algebraic Geometry Seminar, March 2004

• “Higher Dimensional Algebraic Geometry” Conference, Taipei, March 2004

• Western Algebraic Seminar (WAGS), Vancouver, September 2003

• Mathematische Arbeitstagung, Bonn 2003

• “Geometry of Moduli Spaces”, Lille, June 2003

• Euroconference “Vector Bundles on Algebraic Curves”, Porto, July 2003

• Princeton Algebraic Geometry Seminar, April 2003

• “K-theory and Algebraic Cycles”, Miami University, Ohio, April 2003.

• Euroconference “Vector Bundles on Algebraic Curves”, Luminy, July 2002

• Europroj conference “Higher Dimensional Complex Geometry”, Newton Institute, Cambridge, June 2002

• “Classical Algebraic Geometry”, Oberwolfach, May 2002

• Newton Institute, Cambridge, Program on Higher Dimensional Complex Geometry, June 2002.

• Harvard-MIT Algebraic Geometry Seminar, February 2002

• Invited Special Session Talk at the AMS Sectional Meeting Ohio State University, Columbus, September 2001

• Europroj Algebraic Geometry Conference 2001, Norfjordeid, Norway, June 2001 • Europroj Annual Conference 1999, Kibutz Ma’agan, Israel, September 1999

• “Free Resolutions and Linear Systems”, Bayreuth, July 1999

• Harvard-MIT Algebraic Geometry Seminar, November 1998

• European Summer School “Spaces with Singularities”, Cluj (Kolozsvár), July 1998

• Colloquium talks:

2000: Universiteit van Amsterdam 2002: University of Texas at Austin 2003: University of British Columbia, Indiana University at Bloomington, University of California at Irvine, Penn State University, University of Illinois at Urbana-Champaign 2004: Yale University, Rice University 2005: University of Zürich, New Mexico State University, University of San Antonio 2006: University of Georgia, Humboldt Universität zu Berlin, University of Warwick 2007: University of Linnköping (Sweden) 2008: Berlin Mathematical School, Universität Bonn 2009: Ruprecht-Karls-Universität Heidelberg, Universität Hamburg, Université “Joseph Fourier”- Grenoble 2010: Stony Brook, University of Georgia, Universität Freiburg, Seminario Matematico e Fisico di Milano, Johannes Gutenberg Universität Mainz 2012: Leibniz Universität Hannover, Institute of Mathematics of the Romanian Academy 2014: Ludwig-Maximillian Universität München 2015: Babes-Bolyai University, Kolozsvár (Cluj), Berlin Mathematical School, Seminario Matematico e Fisico di Milano 2016: Universität des Saarlandes, Technische Universität Darmstadt, Bar Ilan University 2018: Universität Augsburg 2019: Université Montpellier, Notre Dame University, University of Tokyo 2021: Universität Hamburg 2022: Felix Klein Kolloquium, Universität Leipzig

• Seminar talks:

USA: University of Chicago (2004, 2006, 2019), Harvard University (1999, 2002, 2004, 2010, 2021), University of Pennsylvania (2005, 2006), University of California-Berkeley (2005, 2009), University of California-San Diego (2021), Columbia University (2003, 2021), Johns Hopkins University (2003), Texas A&M University (2005, 2007), Stanford University (2004, 2017, 2020), University of Georgia (2006, 2010), University of Illinois Chicago (2019), University of Texas (2002), University of Illinois Urbana-Champaign (2003, 2004), Rice University (2004), Yale University (2004), Ohio State University (2006), University of Massachusetts at Amherst (2007), University of Michigan (2000, 2002, 2013), SUNY Stony Brook (2010), University of Wisconsin (2021)

Canada: University of British Columbia (2003, 2004)

Germany: Friedrich-Alexander Universität Erlangen-Nürnberg (2008), Universität Duisburg-Essen (2008, 2013), Universität Göttingen (2009), Universität des Saarlandes (2009, 2013), Johannes Guttenberg Universiät Mainz (2010), Universität Bayreuth (2012), Max Planck Institute Bonn (2003, 2006, 2008), Universität Bonn (2015, 2018), May Planck Institute Leipzig (2016, 2018, 2022)

Spain: University of Barcelona (2008, 2015)

Italy: Università di Roma 3 (2008, 2011, 2013, 2014, 2015, 2019, 2022), Università di Genova (2010), Università di Roma 1 Sapienza (2011, 2014, 2015), Università di Roma 2 Tor Vergata (2011, 2015), Università di Milano (2010), Instituto Nazionale di Alta Matematica (2011), Università di Pavia (2013, 2017) Università di Firenze (2015)

England: Cambridge University (2002, 2011), Warwick University (2006), Oxford University (2006), Imperial College London (2016)

France: Université Pierre et Marie Curie Paris 6 (2010, 2017, 2022), Université Joseph Fourier Grenoble (2009), Université de Bourgogne Dijon (2015), Université de Lille (2016), Université Paris 7 (2016)

Hungary: Rényi Institute of Mathematics Budapest (2013, 2016, 2017, 2019, 2020, 2021, 2022)

The Netherlands: Universiteit Utrecht (1999), Universiteit van Amsterdam (1999, 2000, 2012, 2013, 2015), Universiteit Leiden (2006)

Switzerland: ETH Zürich (2011, 2013), Université de Genéve (2013)

Sweden: Mittag-Leffler Institute (2007,2021, 2022), KTH Stockholm (2011, 2012)

Korea: Sogang University (2009), KIAS Seoul (2005, 2009), KAIST Daejon (2007, 2013) Japan: University of Tokyo (2007, 2019) Romania: Institutul de Matematica al Academiei Romane, Bucuresti (2012, 2015) Austria: Institute of Science and Technology Austria (ISTA) (2022) Poland: University of Warsaw (2019)

Languages

English, German, Hungarian and Romanian